# Homophily-Driven Sanitation View for Robust Graph Contrastive Learning

### v2

@@ Abstract
We investigate the adversarial robustness of unsupervised Graph Contrastive Learning (GCL) against structural attacks. First, we provide a comprehensive empirical and theoretical analysis of existing attacks, revealing how and why they downgrade the performance of GCL. Inspired by our analytic results, we present a robust GCL framework that integrates a homophily-driven sanitation view, which can be learned jointly with contrastive learning. A key challenge this poses, however, is the non-differentiable nature of the sanitation objective. To address this challenge, we propose a series of techniques to enable gradient-based end-to-end robust GCL. Moreover, we develop a fully unsupervised hyperparameter tuning method which, unlike prior approaches, does not require knowledge of node labels. We conduct extensive experiments to evaluate the performance of our proposed model, GCHS (Graph Contrastive Learning with Homophily-driven Sanitation View), against two state-of-the-art structural attacks on GCL. Our results demonstrate that GCHS consistently outperforms all state-of-the-art baselines in terms of the quality of generated node embeddings as well as performance on two important downstream tasks [^1] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible..
@@
Graph contrastive learning, Graph sanitation, Adversarial robustness
@@ I Introduction
Graph representation learning [1, 2] has revolutionized the analysis of graph data, which is prevalent across diverse domains. In practice, the scarcity of ground-truth labels has led to a surge in research on unsupervised graph learning approaches. Among these methods, Graph Contrastive Learning (GCL) [3, 4, 5, 6] has emerged as a highly effective unsupervised approach, outperforming other methods on various downstream tasks, including node classification and graph clustering [3, 4, 5].
@@
The label-preserving property of GCL is a key factor behind its main benefits [7]. Specifically, the node embeddings generated by GCL, which incorporate both topological and semantic information, are consistent with node label information even without explicitly querying the node labels. This consistency arises from the homophily [8] assumption of the graph data, which states that nodes tend to connect with “similar” others. This phenomenon is widely observed in real-world graph data such as friendship networks [8], political networks [9], citation networks [10], etc. By contrasting positive and negative samples of nodes that are similar or dissimilar in semantic information, the GCL framework encourages the GNN encoder to learn node embeddings that capture the homophily patterns present in the graph.
@@
However, similar to the vanilla Graph Neural Network (GNN) [11] and its variants, GCL models’ label-preserving property also makes them susceptible to graph structural attacks [12, 13]. In such attacks, the adversary can manipulate the graph topology by adding or deleting edges in the original graph to undermine the quality of node embeddings and subsequently degrade performance on downstream tasks. For instance, a malicious entity can manipulate social ties by connecting with normal accounts to evade graph-based anti-fraud systems [14]. Indeed, existing research has already demonstrated the vulnerability of GCL under structural attacks. In particular, CLGA [13] is an attack method specifically designed against GCL and has demonstrated effective attack performance. Moreover, although Mettack [12] is an attack method designed to attack semi-supervised GNN, extensive experiments [13] demonstrate that it can also successfully attack GCL, at times generating attacks that are more potent than CLGA. Therefore, investigating potential countermeasures against structural attacks is essential to ensuring security and robustness of GCL.
Our main goal is to design the GCL framework that is robust against attacks on the graph structure. While there have been a number of advances in devising countermeasures against structural attacks for GNN models [15, 16, 17, 18], these are primarily designed for semi-supervised settings and intimately depend on having access to node label information. Since GCL is fully unsupervised, such methods are not readily applicable.
Two recent approaches, ARIEL and SPAN, tackle specifically the problem of learning robust GCL models. ARIEL [5] employs an adversarial training approach by introducing an adversarial view into the contrastive learning process to enhance robustness. SPAN [6], on the other hand, modifies the topology augmentation process by maximizing and minimizing the spectrum of the original graph to generate augmentation views. However, these methods have significant limitations. While adversarial training is effective to combat attacks at inference time, their use to counter poisoning attacks on graph structure has the effect of further poisoning the structure of the graph. As a consequence—as we show in the experiments below—ARIEL exhibits limited effectiveness as we increase attack strength. Moreover, adversarial training is tied to the specific mechanism presumed to poisoned the graph structure, making ARIEL vulnerable whenever actual attacks make use of different poisoning techniques. As for SPAN, minimizing the spectrum of graph Laplacian will naturally add inter-class edges (similar to adversarial noise) to the input graph, inheriting similar limitations as adversarial training. In addition, computing spectral changes is quite time-consuming due to the involvement of eigenvalue decomposition of the graph Laplacian ($\mathcal{O}(N^{3})$ where $N$ is the node number). Indeed we observe that SPAN demonstrates better robustness than ARIEL because of an augmentation view generated by maximizing the spectrum of graph Laplacian, which can prune the inter-class edges, resulting in a “cleaner” graph. This observation suggests a more direct approach for sanitizing the graph that is focused specifically on recovering homophily. Finally, both SPAN and ARIEL tune the vital hyperparameters based on the performance on downstream tasks, requiring access to node labels. This takes them partly outside the unsupervised setting that motivates GCL.
We propose a robust GCL framework that integrates a homophily-driven learnable sanitation view that can effectively prune potential malicious edges in the poisoned graph during training. Of course, the central challenge is how to effectively design such a sanitation view. To address this challenge, we first investigate the vulnerability of GCL under state of the art structural attacks. This investigation yields two crucial insights. Firstly, extensive experiments demonstrate that the graph structural attacks will break an important graph property: graph homophily [8] (i.e., the tendency of similar nodes to be connected). Secondly, we theoretically prove that structural attacks will adversely affect the mutual information of the graph and its representations from the information-theoretic perspective. Based on these results, we propose to learn the sanitation view under the guidance of both graph homophily and mutual information.
There are several technical challenges in effectively training our approach in a fully end-to-end manner. Firstly, measuring graph homophily typically requires information about node labels, which is missing in our unsupervised setting. Secondly, augmenting topology changes to graph data is a complex discrete optimization problem not directly amenable to gradient-based learning techniques. Thirdly, all prior approaches fine-tune hyperparameters by utilizing supervised information about node labels; while this is infeasible in a fully unsupervised setting, it is not evident how to avoid it.
We propose a series of techniques to address the above challenges to enable full end-to-end gradient-based robust GCL. Specifically, we use a random vector following the Bernoulli distribution to control the generation of the sanitation view, where the parameters of the distribution are optimized via learning. During the gradient update phase, we employ the Gumbel-Softmax re-parametrization trick [19] to approximate the Bernoulli distribution and use the straight-through estimator [20] to address non-differentiability. The entire framework is then optimized in an end-to-end manner, training the learnable sanitation view generator and GNN encoder simultaneously. Our approach is therefore considerably more time-efficient than the two-stage GCLs such as SPAN. Additionally, to measure graph homophily without the knowledge of node labels, we use the raw node attribute matrix as an alternative. Moreover, we implement the early stopping technique based on a pseudo normalized cut loss to determine the best choice of the vital hyperparameter $\eta$. We refer to our proposed robust model as GCHS (Graph Contrastive Learning with Homophily-Driven Sanitation View).
We conduct extensive experiments to evaluate GCHS. We compare GCHS with seven strong baselines and report robustness results against two representative state-of-the-art structural attacks over six widely-used graph datasets. Our evaluation considers two important downstream tasks for GCL: node classification and graph clustering. Our results demonstrate that GCHS consistently outperforms all baselines in nearly all cases, often by a large margin.
@@
  • We propose a novel GCL framework that integrates a learnable sanitation view, which significantly improves the adversarial robustness of unsupervised graph contrastive learning under structural attacks.
  • We propose a series of techniques to enable effective end-to-end gradient-based training of the learnable sanitation view generator and GCL.
  • We propose a pseudo normalized cut loss to conveniently fine-tune the vital hyperparameter $\eta$ without altering the GCL’s model structure, and thus entirely avoid the need for node labels, unlike past approaches for GCL models.
  • Extensive experiments demonstrate the effectiveness of our proposed robust GCL framework under different attack settings.
@@ II-A Graph Structural Attack
Graph-based machine learning models have been shown to be vulnerable to structural attacks. Nettack [21] was the first targeted attack on GNNs that manipulated both on node attribute matrix and adjacency matrix. Mettack [12] formulated the global structural poisoning attacks on GNNs as a bi-level optimization problem and leveraged a meta-learning framework to solve it. BinarizedAttack [14] simplified graph poisoning attacks against the graph-based anomaly detection to a one-level optimization problem and optimize it by mimicking the training of the binary neural network. HRAT [22] proposed a heuristic optimization model integrated with reinforcement learning to optimize the structural attacks against Android Malware Detection. CLGA [13] deployed GCA [4] as the surrogate model and formulated graph structural attacks as a bi-level optimization problem to degenerate the performance of GCL models through poisoning. In this paper, we choose two typical graph structural attacks Mettack and CLGA as our graph attackers to simulate real-world attacking scenarios.
@@ II-B Graph Contrastive Learning
Contrastive learning is a widely-used deep learning model that originated from SimCLR [23], a simple self-supervised contrastive framework for learning useful representations for image data. SimCLR introduces the infoNCE loss to approximate the mutual information between latent representations from different augmentation views, such as randomly cropping and resizing, rotation, Gaussian blurring and color distortion. GRACE [3] extended the SimCLR framework to graph data by generating augmentation views through random link removal and feature masking. Later, GCA [4] prevents the removal of the important links during the stochastic augmentations by designing several link removal mechanisms based on degree centrality, eigenvector centrality and PageRank centrality. ARIEL [5] introduced an adversarial view via PGD attack and an information regularization penalty to stabilize training. SPAN [6] explored spectrum invariance during augmentation views and generated augmentation views by maximizing and minimizing the spectral change. MVGRL [24] utilized a local-global contrastive loss to measure the agreement between the neighbor nodes and graph diffusion. DGI [25] generated augmentation views by corrupting the original graph and contrasting the node embeddings between the original graph and the corrupted graphs. Finally, BGRL [26] introduced bootstrapped graph latent training by predicting alternative augmentations of the input without contrasting with negative samples.
@@ III Preliminaries
In this section, we will provide a brief introduction to the prerequisite knowledge and notations to our work.
@@ III-B Graph Contrastive Learning Framework
GCL is a powerful graph representation learning method that seeks to maximize the similarity between positive pairs (same node with different views) and enlarge the difference between negative pairs (different nodes within the same view and cross different views). More specifically, the training of the GCL model has two stages. Firstly, it generates two augmentation views $G_{1}$ and $G_{2}$ based on the input graph $G$ via pre-defined augmentation operations such as random link removal and feature masking. Then, Two graph samples generated from augmentation views are fed into a shared GNN encoder to produce node embeddings. Finally, a contrastive loss (infoNCE loss) based on the node embeddings is optimized and update the parameters of the GNN encoder. The infoNCE loss is formulated as:
@@
where $u_{i}$ is the $i$-th node for $G_{1}$ and $v_{i}$ is the $i$-th node for $G_{2}$, $\tau$ is a temperature hyperparameter. Usually, the similarity function $\rho(\cdot)$ is defined as:
@@
which represents the cosine similarity between the projected embeddings of $u_{i}$ and $v_{i}$, $g(\cdot)$ is the projection head to enhance the expressive power of the graph representation learning, $\mathbf{H}$ is the node embeddings to be detailed later. In this paper, we use the two-layered graph convolution [11] as the shared GNN encoder for both two views to get the node embeddings $\mathbf{H}_{1}=f_{\theta}(G_{1})$ and $\mathbf{H}_{2}=f_{\theta}(G_{2})$ respectively. Specifically, $G_{1}=\{\mathbf{A}_{1},\mathbf{X}_{1}\}$ and $G_{2}=\{\mathbf{A}_{2},\mathbf{X}_{2}\}$; then we have
@@
Then, the optimized node embeddings $\mathbf{H}^{*}=f_{\theta^{*}}(\mathbf{A},\mathbf{X})$ act as the inputs to the logistic regression for semi-supervised node classification.
III-C Normalized Cut
Normalized cut loss [27, 28] is an unsupervised loss function that counts the fraction of the inter-class links for a given graph partition result. Thus, a good graph partition should obtain a lower normalized cut loss, i.e., the majority of the links connect two nodes with the same class. The normalized cut loss is defined as:
$\begin{split}&\mathcal{L}_{nc}=\frac{1}{K}\operatorname{Tr}((\mathbf{C}^{\top}\mathbf{LC})\oslash(\mathbf{C}^{\top}\mathbf{DC})),\ \mathbf{L}=\mathbf{I}-\hat{\mathbf{A}}.\end{split}$ (2)
[…]
IV Problem Statement
In this section, we introduce the threat model and formally describe the problem of defense of GCL.

### v3

# Robust Graph Contrastive Learning with Information Restoration

@@ Abstract
The graph contrastive learning (GCL) framework has gained remarkable achievements in graph representation learning. However, similar to graph neural networks (GNNs), GCL models are susceptible to graph structural attacks. As an unsupervised method, GCL faces greater challenges in defending against adversarial attacks. Furthermore, there has been limited research on enhancing the robustness of GCL. To thoroughly explore the failure of GCL on the poisoned graphs, we investigate the detrimental effects of graph structural attacks against the GCL framework. We discover that, in addition to the conventional observation that graph structural attacks tend to connect dissimilar node pairs, these attacks also diminish the mutual information between the graph and its representations from an information-theoretical perspective, which is the cornerstone of the high-quality node embeddings for GCL. Motivated by this theoretical insight, we propose a robust graph contrastive learning framework with a learnable sanitation view that endeavors to sanitize the augmented graphs by restoring the diminished mutual information caused by the structural attacks. Additionally, we design a fully unsupervised tuning strategy to tune the hyperparameters without accessing the label information, which strictly coincides with the defender’s knowledge. Extensive experiments demonstrate the effectiveness and efficiency of our proposed method compared to competitive baselines.
@@
Robust Graph Contrastive Learning, Graph Representation Learning, Adversarial Robustness
@@ I Introduction
Graph representation learning [1, 2] has revolutionized the analysis of graph data, which is prevalent across diverse domains. In practice, the scarcity of ground-truth labels has led to a surge in research on unsupervised graph learning approaches. Among these methods, Graph Contrastive Learning (GCL) [3, 4, 5, 6, 7] has emerged as a highly effective unsupervised approach, outperforming other methods on various downstream tasks, including semi-supervised node classification [8].
@@
The label-preserving property of GCL is a key factor behind its main benefits [9]. Specifically, the node embeddings generated by GCL, which incorporates both topological and semantic information, are consistent with node label information even without explicitly querying the node labels. This consistency arises from the homophily [10, 11, 12, 13] assumption of the graph data, which states that nodes tend to connect with “similar” others. This phenomenon is widely observed in real-world graph data such as friendship networks [10], political networks [14], citation networks [15], etc. By contrasting positive and negative samples of nodes that are similar or dissimilar in semantic information, the GCL framework encourages the GNN encoder to learn node embeddings that capture the homophily patterns present in the graph.
@@
However, similar to the vanilla Graph Neural Network (GNN) [8, 16] and its variants, GCL models are also susceptible to graph structural attacks [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. In such attacks, the adversary can manipulate the graph topology by adding or deleting edges in the original graph to undermine the quality of node embeddings. For instance, a malicious entity can manipulate social ties by connecting with normal accounts to evade graph-based anti-fraud systems [30]. Indeed, existing research has already demonstrated the vulnerability of GCL under structural attacks. In particular, CLGA [19] is an attack method specifically designed against GCL and has demonstrated effective attack performance. Moreover, although Mettack [18] is an attack method designed to attack semi-supervised GNN, extensive experiments [19] demonstrate that it can also successfully attack GCL, at times generating attacks that are more potent than CLGA. Therefore, investigating potential countermeasures is essential to ensuring the security and robustness of GCL.
Our main goal is to refine the GCL framework so that it is robust against graph structural attacks. While there have been a series of advances in devising countermeasures against structural attacks for GNN models [31, 32, 33, 34, 35, 13, 36, 37, 38, 39], they are primarily designed for semi-supervised settings and intimately require access to node label information. Since GCL is fully unsupervised, such methods are not readily applicable. Two recent approaches, ARIEL [5] and SPAN [6], specifically address the challenge of learning robust GCL models. ARIEL [5] employs an adversarial training approach by introducing an adversarial view into the contrastive learning process to improve robustness. SPAN [6], on the other hand, modifies the topology augmentation process by maximizing and minimizing the spectrum of the original graph to generate augmentation views. However, the major drawback with these methods is that their strategy to mitigate graph structural attacks has the effect of further poisoning the graph’s structure. For instance, the adversarial view of ARIEL injects adversarial noises into both the feature and topology spaces of the original poisoned graph. Nevertheless, the adversarial noises could still have persistent negative impact on graph representation learning, leading to limited defense performance. The same problem also exists in one of the augmented views of SPAN whose goal is to minimize the spectrum of the graph Laplacian.
To avoid the above problem, in this paper, we take a quite different angle to achieve adversarial robustness, by restoring the graph’s properties contaminated by adversarial attacks. Intuitively, refining the poisoned graph is more effective than adversarial training since the latter cannot omit the harmful effects of the existing adversarial noises. Besides, the observation that SPAN achieves better robustness than ARIEL is because of an augmented view generated by maximizing the spectrum of graph Laplacian, which can restore the diminished graph homophily [10] and thus lead to a “cleaner” graph. Therefore, restoring the graph property is a promising and effective way to improve GCL robustness against attacks.
To this end, we begin by assessing the vulnerability of the GCL framework to graph structural attacks. Our focus is on identifying important graph properties that are significantly changed during the attack. Apart from the common observation that structural attacks will typically reduce graph homophily, we discover that these attacks will essentially degrade the mutual information estimation between the graph and its representations (refer to Sec. IV-B). This sheds light on how attacks affect the graph learning tasks from an information-theoretic perspective. Specifically, the mutual information between the graph and its representations serves as a crucial cornerstone for generating high-quality node representations within the GCL framework. The graph attacker undermines the quality of the GCL’s node embeddings by reducing the alignment between the graph’s information (including semantic and topology information) and the low-dimensional representations. Consequently, the poisoned GCL’s node embeddings fail to capture the “true” information of the graph data. This critical observation underscores the significance of restoring mutual information to effectively defend against such attacks.
Furthermore, GCL models are unsupervised approaches, meaning that the GCL’s node embeddings are pre-trained without accessing the node labels. However, the current GCL models [4, 5, 6, 40, 41] often rely on the validation label set to tune the hyperparameters of the GCL framework, which contradicts the principles of unsupervised learning. Therefore, training robust GCL models in a fully unsupervised setting in an adversarial environment encounters two major challenges: 1) How to restore the diminished mutual information caused by the noisy data during the training of the GCL framework? 2) How to determine the vital hyperparameters of the GCL framework in a fully unsupervised setting?
To address the challenges, we propose a robust GCL framework termed GCIR that endeavors to restore diminished mutual information after structural attacks, thus achieving adversarial robustness. Specifically, GCIR novelty integrates a learnable sanitation view into the framework and contrasts it with an augmented view (through feature masking or link dropping [3, 4]) to gradually restore the diminished mutual information. The sanitation view is trained jointly with the GNN encoder in an end-to-end manner. To train the model in a fully unsupervised manner, we design an early stopping strategy based on a pseudo-normalized cut [42] metric without requiring the label’s information for the downstream tasks to determine the best choice of the vital hyperparameter.
@@
  • We provide the vulnerability analysis of the GCL framework under graph structural attacks from an information theoretical perspective and verify that graph structural attacks will degrade the mutual information estimation between the graph and its representations. This finding advances our understanding of attacks and provides valuable insights for developing defense approaches.
  • We propose a novel robust GCL framework, termed GCIR, that achieves adversarial robustness by introducing a learnable sanitation view to restore the diminished mutual information during training. Moreover, we design an unsupervised tuning strategy to determine the best choice of the vital hyperparameter without label information.
  • Extensive experiments demonstrate the effectiveness of our proposed robust GCL framework across various attack scenarios.
@@ II-A Graph Structural Attack
Graph-based machine learning models have been shown to be vulnerable to structural attacks. Mettack [18] formulated the global structural poisoning attacks on GNNs as a bi-level optimization problem and leveraged a meta-learning framework to solve it. CLGA [19] deployed GCA [4] as the surrogate model and formulated graph structural attacks as a bi-level optimization problem to degenerate the performance of GCL models through poisoning. There are other structural attacks against graph-based machine learning models [30, 20, 21, 43] such as graph anomaly detection, malware detection, random walks, etc.
@@ II-B Graph Contrastive Learning
Contrastive learning is a widely used deep learning model that originated from SimCLR [44], a simple self-supervised contrastive framework for learning useful representations for image data. GRACE [3] extended the SimCLR framework to graph data by generating augmentation views through random link removal and feature masking. Later, GCA [4] prevents the removal of the important links during the stochastic augmentations by designing several link removal mechanisms based on degree centrality, etc. ARIEL [5] introduced an adversarial view via a PGD attack for robust training. SPAN [6] explored spectrum invariance during augmentation views and generated augmentation views by maximizing and minimizing the spectral change. MVGRL [40] utilized a local-global contrastive loss to measure the agreement between the neighbor nodes and graph diffusion. DGI [41] generated augmentation views by corrupting the original graph and contrasting the node embeddings between the original graph and the corrupted graphs. BGRL [45] introduced bootstrapped graph latent training by predicting alternative augmentations of the input without contrasting with negative samples. SPAGCL [46] integrated the node similarity-preserving view and the adversarial view to enhance the adversarial robustness of GCL model. PiGCL [47] endeavored to mitigate the implicit conflicts from the negative pairs caused by the InfoNCE loss of GCL to boost its robust performance. GPS [48] deployed a special pooling mechanism to enhance the adversarial robustness of GCL on the graph classification problem.
However, the above-mentioned (robust)-GCL methods contain their unique limitations to some extent. For example, the adversarial view of ARIEL and SPAGCL injects adversarial noises into the feature and topology space of the graph data during training, which will still hinder the robust learning of GCL. The view that minimizes the spectral norm will introduce more inter-class links and produce similar detrimental effects on the graph representation learning. On the other hand, although PiGCL endeavors to improve the robustness of GCL by refining the gradients of the negative pairs, it still cannot essentially erase the adverse impact from the topology space of the graph data. Lastly, the adversarial pooling of GPS is specifically designed to boost the adversarial robustness of GCL on the graph classification task, which falls outside the scope of our topic (robustness of node classification). It is worth noting that unlike other baselines, our work focuses on sanitizing the detrimental effects of the adversarial attacks on the input space, which is supervised by a specially crafted sanitation view of the GCL framework from an information-theoretical perspective.
@@ III Preliminaries
In this section, we will provide a brief introduction to the prerequisite knowledge and notations for our work.
@@ III-B Graph Contrastive Learning Framework
GCL is a powerful graph representation learning framework that seeks to obtain high-quality node embeddings via maximizing the intractable mutual information between the graph data and its low-dimensional representations, i.e.,
$I(G,\mathbf{Z})=\int_{z}\int_{g}p(g,z)\log(\frac{p(g,z)}{p(g)p(z)})dgdz.$
In order to achieve this goal, a common way is to maximize the similarity between positive pairs (same node with different views) and enlarge the difference between negative pairs (different nodes within the same view and cross different views). In general, the training of the GCL model has two stages. Firstly, it generates two augmentation views $G_{1}$ and $G_{2}$ based on the input graph $G$ via random link removal and feature masking. Then, Two graph samples generated from augmentation views are fed into a shared GNN encoder to produce node embeddings. Finally, a contrastive loss (InfoNCE loss) based on the node embeddings is optimized. The InfoNCE loss is formulated as:
@@
where $u_{i}$ is the $i$-th node for $G_{1}$, $v_{i}$ is the $i$-th node for $G_{2}$, and $\tau$ is a temperature hyperparameter. Usually, the similarity function $\rho(\cdot)$ is defined as:
@@
which represents the cosine similarity between the projected embeddings of $u_{i}$ and $v_{i}$, $g(\cdot)$ is the projection head to enhance the expressive power of the graph representation learning, $\mathbf{H}$ is the node embeddings to be detailed later. We use the two-layered graph convolution [8] as the shared GNN encoder for both two views to get the node embeddings $\mathbf{H}_{1}=f_{\theta}(G_{1})$ and $\mathbf{H}_{2}=f_{\theta}(G_{2})$ respectively. Specifically, $G_{1}=\{\mathbf{A}_{1},\mathbf{X}_{1}\}$ and $G_{2}=\{\mathbf{A}_{2},\mathbf{X}_{2}\}$; then we have
@@
Then, the optimized node embeddings $\mathbf{H}^{*}=f_{\theta^{*}}(\mathbf{A},\mathbf{X})$ serve as the input to the logistic regression for semi-supervised node classification.
IV Vulnerability Analysis of GCL
In this section, we present a detailed theoretical and empirical analysis of how existing attacks influence GCL, focusing on key graph properties. These explorations of the vulnerabilities of GCL provide strong insights for our design of the robust GCL framework. Without loss of generality, we choose two graph structural attacks, i.e., Mettack [18] and CLGA [19] as our graph attackers to simulate real-world attacking scenarios. Specifically, Mettack is the most representative structural attack against GNNs, as other ones [21] share similar attack loss with Mettack. Meanwhile, CLGA is the up-to-date representative attack method against the GCL framework. Therefore, exploring the underlying principles of Mettack and CLGA can sufficiently reveal the vulnerability of the GCL framework under structural attacks.